OpenStudy (anonymous):

Is the sequence geometric? 2, 6, 18, 54,... A. yes B. no

OpenStudy (anonymous):

yes since 6/2 = 18/6 = 54/18 = 3

OpenStudy (anonymous):

A You have to check if the common ration is equal between every an a(n) and a(n+1) 6/2 = 18/3 = 54/18 = 3 = common ratio